OpenStudy (anonymous):

find the derivative of f(x)=3x^2-5x using f'(x)= lim y->x(f(y)-f(x))/y-x

OpenStudy (anonymous):

where are you stuck??

OpenStudy (anonymous):

Im not sure how to plug it in. I should put it in for x would I? But what do I do with Y?

OpenStudy (anonymous):

\[ \lim_{y\to x}\frac{f(y)-f(x)}{y-x} = \lim_{y\to x}\frac{3y^2-5y-(3x^2-5x)}{y-x} \]

OpenStudy (anonymous):

Can you do it from there?

OpenStudy (anonymous):

Not really. Do i need to take out an x and y?

OpenStudy (anonymous):

You need to simplify a bit.

OpenStudy (anonymous):

\[ \lim_{y\to x}\frac{3y^2-5y-(3x^2-5x)}{y-x} = \lim_{y\to x}\frac{(3y^2-3x^2)-(5y-5x)}{y-x} \]

OpenStudy (anonymous):

Getting it yet?

OpenStudy (anonymous):

couldn't i do it like this \[y(3y-5)-x(3x-5)/ y-x\]

OpenStudy (anonymous):

\[ \lim_{y\to x}\frac{(3y^2-3x^2)-(5y-5x)}{y-x} = \lim_{y\to x}\frac{3(y^2-x^2)-5(y-x)}{y-x} \]

OpenStudy (anonymous):

\[ \lim_{y\to x}\frac{3(y^2-x^2)-5(y-x)}{y-x} = \lim_{y\to x}\frac{3(y-x)(y+x)-5(y-x)}{y-x} \]

OpenStudy (anonymous):

\[ \lim_{y\to x}3(y+x)-5 \]

OpenStudy (anonymous):

@ns585 Now do you get it? I had to do difference of squares.

OpenStudy (anonymous):

yes i see what you did but don't I have to do something else also or is that the answer?

OpenStudy (anonymous):

First of all you need to apply the Sum and Difference rule. that is\[\frac{ d }{ dx } (f(x)-g(x))= \frac{ d}{ dx } (f(x))- \frac{ d }{ dx } (g(x))\] In your case is\[f(x)=3x^2-5x\] then we just follow the rule\[f(x)= \frac{ d }{ dx }(3x^2)-\frac{ d }{ dx}(5x)\] use the power rule for 3x^2 that should be\[n x^n-1\] at the end it would be\[f(x)=6x-5\]

OpenStudy (anonymous):

sorry my bad the power rule is\[n x^{n-1}\]

OpenStudy (anonymous):

\[ \lim_{y\to x}3(y+x)-5=3(x+x)-5=6x-5 \]